OpenStudy (anonymous):

What is the value of expression 2/3 + 1/6

OpenStudy (anonymous):

what do you think is the answer?

OpenStudy (anonymous):

I don't know thats why i'm asking I new to this

OpenStudy (anonymous):

First find the greatest common factor of the two denominator 3 and 6. What is the greatest common denominator of those two numbers?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

oh ha a little off

OpenStudy (anonymous):

\(\huge \frac{2}{3} + \frac{1}{6} \) The GCD of 3 and 6 is 6

OpenStudy (anonymous):

oh because they all go into each other

OpenStudy (anonymous):

\(\huge \frac{2}{3} + \frac{1}{6} \) The GCD of 3 and 6 is 6 Now we have a common denominator, which is 6 \(\huge \frac{2}{3} \frac{?}{6} + \frac{1}{6} \frac{?}{6}\) \(\huge \frac{2}{3} \frac{4}{6} + \frac{1}{6} \frac{1}{6}\) \(\huge \frac{4}{6} + \frac{1}{6} = ?\)

OpenStudy (anonymous):

5/6

OpenStudy (anonymous):

Correct!!

OpenStudy (anonymous):

i thought it was four

OpenStudy (anonymous):

ok wb this one If QRST below is a rectangle, find x if angle T is equal to Imported Asset .